Glucose Scan Incorporated (GSI) currently sells its latest glucose monitor, the Glucoscan 3000, to diabetic patients for $129. GSI plans on lowering their price next year to $99 per unit. The cost of goods sold for each Glucoscan unit is $50, and GSI expects to sell 100,000 units over the next year.

(1) Suppose that if GSI drops the price on the Glucoscan 3000 immediately, it can increase sales over the next year by 35% to 135,000 units. What is the incremental impact of this price drop on the firms EBIT? (Hint: EBIT=Sales-COGS)

(2) Suppose that if GSI drops the price on the Glucoscan 3000 immediately, it can increase sales over the next year by 35% to 135,000 units. Also suppose that for each Glucoscan monitor sold, GSI expects additional sales of $100 per year on glucose testing strips and these strips have a gross profit margin of 70%. Considering the increase in the sale of testing strips, what is the incremental impact of this price drop on the firms EBIT? (Hint: EBIT=Sales-COGS)
Answer #1

rate positively ..

Ans a)
Computation of original EBIT
sales         12,900,000
129*100000
Less Cost           5,000,000
50*100000
A EBIT           7,900,000
Revised EBIT
sales 13365000
99*135000
Less Cost 6750000
50*135000
B EBIT           6,615,000
C=B-A Incremental impact =         (1,285,000)
EBIT will decline by 1285000
Ans b)
A Base EBIT=           7,900,000
Testing strips incremental EBIT
sales 3500000
100*35000
EBIT @70% 2450000
EBIT =           2,450,000
B Revised total EBIT=           9,065,000
4050000+6615000
C=B-A Incremental impact           1,165,000